Assume that you are using qPCR to determine the amount of DNA in a given sample. Sample 1, from which you know that it contained an amount of 1 pg, required 15 PCR cycles to reach a specific level of PCR product. Your sample 2 required only 5 PCR cycles to reach the same level. Sample 3 required 18 PCR cycles to come to the same level of PCR product. How much DNA did samples 2 and 3 contain?

RT-qPCR can be used to quantify the amount of specific RNA molecules in a sample. However, PCR is able to amplify only DNA. How is it possible that RT-qPCR can quantify the amount of RNA?
Answer #1

Amount of DNA in sample 1= 1pg

After 15 cycles the no of DNA double stranded pieces= 215=32,768

After 15 cycles amount of DNA=32768 pg

For sample 2 cycle required=5

For sample 2 after 5 cycles the no of DNA double stranded pieces=25=32

so amount of DNA in sample 2 at start point=32768/32=1024 pg

For sample 3, Cycles required=18

For sample 3 after 18 cycles the no of DNA double stranded pieces=262144

so amount of DNA in sample 3 at start point=32768/262144=0.125 pg

The PCR is designed in such a way that it can only amplify the target gene sequence by using the specific primers and no dye is used in the normal PCR. While in case of qPCR the fluorescent make qPCR more quantitative.